OpenStudy (anonymous):

how would you calculate the population variance of a data set given a normal distribution with 45 data points, a mean of 15, median of 15 and sample variance of 7.5?

OpenStudy (dumbcow):

the difference between sample variance and pop variance is you divide by N-1 and N respectively Let X be sum of squared deviations \[\frac{X}{N-1} = 7.5\] where N = 45 \[X = (7.5)(44)\] so pop variance is X/N \[\frac{X}{45} = \frac{(7.5)(44)}{45}\]

OpenStudy (anonymous):

Thank you, that makes sense now. I was trying to make it much more difficult than it actually is.